OpenStudy (bluestar70556):

Let f be defined as follows, where a CANNOT equal 0 f(x) = (x^3 - a^3)/(x - a) when x does not equal a f(x) = 3a^2 when x DOES equal a Which of the following are true about f? I. Lim(x -> a) f(x) exists II. f(a) exists III. f(x) is continuous at x = a Note: It is possible for one, none, or all to be correct Somebody please help me??? I'm so confused

OpenStudy (kittiwitti1):

when \(x\ne a\),\[f(x)=\frac{(x^{3}-a^{3})}{(x-a)}\]when \(x=a\),\[f(x)=3a^{2}\]

OpenStudy (kittiwitti1):

This is what you are given, yes?

OpenStudy (bluestar70556):

Yes that is correct (The equations aren't loading on my end but from what I can read that is correct)

OpenStudy (kittiwitti1):

Oh, I'm sorry, here's a screenshot of it: http://prntscr.com/d07gqt

OpenStudy (bluestar70556):

Yep that's correct!

OpenStudy (bluestar70556):

Hang on, I'll be right back, I'm going to log out and log back in, hopefully it will help this glitching out, I'll be back in a second!

OpenStudy (kittiwitti1):

Okay, well... I've just started learning calculus; so PLEASE don't take my words as solid proof, but... this is what I think. (i) cannot be true as the limit at \(a\) is different from both sides. Regarding one-sided functions:\[\lim_{x\rightarrow a^{+}}\ne\lim_{x\rightarrow a^{-}}\] One side would be approaching from \(+\infty\) and the other side would be approaching from \(-\infty\). (ii) is true as there is a point a \(x=a\), it's stated in the problem. (iii) cannot be true unless (i) is true, as continuity means that the limit at \(x=a\) needs to exist (\(I\) \(THINK\)).

OpenStudy (bluestar70556):

That's what I thought too! Since I cannot be true, III cannot be true either but II is possible

OpenStudy (kittiwitti1):

Yes, well ... I would get a second opinion on this, because I'm not quite sure if I have a full grasp on the concepts yet. Also, \(\text{Oakheart says hi.}\) 🐈

OpenStudy (bluestar70556):

LOL! Hello Oakheart my passed love XD I'm glad someone else got the name, and I'll do some more research, but that helped a lot, thanks so much!

OpenStudy (kittiwitti1):

Alright :-) no problem! Good luck ♣ (And... spoiler alert (?) but Brambleclaw becomes Clan Leader haha)

OpenStudy (bluestar70556):

(Oh oops, I guess so XD)) Oh nu! ;-;

OpenStudy (kittiwitti1):

idk how far you are but yeah ! GOOD LUCK ♣ @Directrix @agent0smith help please lol

OpenStudy (agent0smith):

For part I, you need to factor it to check the left and right hand limits. Use a difference of two cubes (google it) \[\large f(x)=\frac{x^{3}-a^{3}}{x-a}\]

OpenStudy (bluestar70556):

Wait a minute if that is factored that gives me a^2 + ax + x^2 and then when I substitute a for x I get a^2 + aa + a^2, which would equal to 3a^2

OpenStudy (agent0smith):

Yes, which means what...? But slow down and show work. You need to find the value of the left and right limits first\[\large \lim_{x \rightarrow a^-}\frac{(x-a)(x^2 +ax + a^2)}{x-a}\]\[\large \lim_{x \rightarrow a^+}\frac{(x-a)(x^2 +ax + a^2)}{x-a}\]are these two going to be equal?

OpenStudy (bluestar70556):

I'm not 100% sure only because I'm not sure how to calculate that and I can't see the equations (my computer is only showing the code) I'm guessing they are going to be equal. I DO know that if I and II are equal to each other that means that they are continuous, so that would mean III is correct as well.

OpenStudy (bluestar70556):

Well, Looking at the equations I would say it's equal.. I'm looking online now to see how to prove this analytically

OpenStudy (agent0smith):

All you have to do is cancel off common factors, and then it's clear they're equal.

OpenStudy (bluestar70556):

So just factor off the (x-a) and then be left with (x^2 + ax + a^2)

OpenStudy (agent0smith):

Yep.

OpenStudy (bluestar70556):

Okay, so In reality all of them would be true, correct?

OpenStudy (bluestar70556):

Because 1 and 2 equal the same thing, so 3 is automatically true.

OpenStudy (bluestar70556):

Also thanks for your help XD

OpenStudy (agent0smith):

Yes, since the limit at a, is equal to f(a), it's continuous.

OpenStudy (bluestar70556):

Perfect, Thanks again for your help.

OpenStudy (agent0smith):

Welcome.